Electrochemical (or voltaic) cell notes AP chemistry Write the re Identify the elements being oxidized and being reduced Describe what you would observe if that reaction were set up like this Where are the electrons being transferred from the zinc atoms to the copper ions? coppertin nitrate soluhon Write the two half reactions that are taking place in this reaction oxidation reduction Lets separate the two half reactions physically. Label how we could do this. wre oses mass gains mass An electrochemical cell of the reaction
media%2F5dd%2F5dd7ff8c-e66c-4916-817e-f2
Answer #1

Ans. Cu(NO3)2 + Zn = Cu(s) + Zn(NO3)2

The element Copper, Cu is being reduced from +2 to +0.

The element Zn gets oxidized from +0 to +2.

This cell is known as Galvanic cell in which Zn metal gets dissoved and form Zn2+ from Zn electrode and Cu2+ gets deposited to Cathode as Cu. and form electrochemical cell and produce electricity.

Oxidation takes place in anode and reduction take place in cathode.

Anode reaction:- Zn(s) = Zn2+ + 2e

Cathode reaction:- Cu2+ + 2e = Cu(s)